User Tools

Site Tools


Problems starting as CA01 had Hw problems. Eventually disabled ca01. Setup took almost 2 hours!!

Calibrated and on-source in the schedule at 23:50 UT.

Tied C2345. Reference CA04.

Disko and good data at 00:01 UT (Had raid access problems)

Persistent problem with DAS2. Error is shown for clock in IFP1. If monitoring, DAS attempts to sync continuously! fauto and fringe checking show that Ch6 has large DC offset and stats only about 15. Power-cycled DAS 10:49-10:50:30. Problem fixed for about 5 minutes but fault came back. Have to live with this for this run.

Tsys logging on CACOR started at 1:47UT. Values before this time can be extracted from MONICA database.

lbaops/lbaaug2007/v234batlog.txt · Last modified: 2015/12/18 16:38 by